# and Semantics for Precise Image Editing

### v1

@@ Abstract
Figure 1: FlashEdit produces superior visual results for text-guided image editing, addressing background instability and semantic entanglement with an over $150\times$ speedup against DDIM (Song et al. (2020b)) + P2P (Hertz et al. (2022)).
@@ 1 Introduction
Text-guided image editing with diffusion models (Brooks et al. (2023),Dong et al. (2023)) has demonstrated remarkable capabilities, allowing users to perform complex semantic modifications with high fidelity. The standard methodology is built upon a two-stage inversion-denoising pipeline: an initial inversion process maps a source image to its corresponding noise latent, which is then progressively denoised to generate the edited output according to a target prompt (Ju et al. (2023),Cao et al. (2023)). The objective is to achieve high fidelity in both content preservation and target prompt alignment, which often necessitates a computationally intensive, multi-step process.
Recent research has pursued several distinct strategies to improve accuracy and speed. To tackle the latency of the multi-step denoising process, methods based on model distillation have been proposed to enable editing in a faster way (Deutch et al. (2024)). These approaches must carefully address challenges such as mismatched noise statistics and insufficient editing strength that arise when adapting multi-step frameworks to fast samplers (Mokady et al. (2023b),Miyake et al. (2025)). To improve edit precision and prevent semantic leakage into the background, another category of work modifies the model’s internal mechanisms, primarily by re-weighting or replacing attention maps to ensure the edit is spatially constrained (Fang et al. (2024); Xu et al. (2024)). Recognizing that the final edit quality is highly dependent on the starting point, other approaches focus on refining the inversion technique itself (Ju et al. (2023)). These methods aim to find a more accurate initial latent vector, with recent insights revealing that separating the objectives of content preservation and edit fidelity can yield significant performance gains and speedups (Wang et al. (2025b)).
@@
  • At the temporal level, we introduce the One-Step Inversion-and-Editing (OSIE) pipeline and its ”Anchor-and-Refine” training strategy, which for the first time enables high-fidelity inversion for one-step diffusion models.
  • At the spatial level, we propose Background Shield (BG-Shield), a structural intervention in self-attention that uses memory caching and selective core querying to enforce pixel-perfect background preservation, ensuring the structural integrity of the edit.
@@ 2.1 Diffusion Models
Recent advances in image synthesis have been largely driven by diffusion models (Peebles & Xie (2023),Kulikov et al. (2024)), which have become a leading paradigm for generating high-fidelity images from text. The core mechanism involves an iterative denoising process that progressively refines a random noise vector into a coherent image conditioned on a text prompt. A landmark contribution in this area is Stable Diffusion  (Rombach et al. (2021)), a Latent Diffusion Model (LDM) (Rombach et al. (2022)) that performs the computationally intensive denoising process in a lower-dimensional latent space, making the technology widely accessible. Parallel to this, alternative frameworks have emerged, such as Flow Matching models like Flux (Labs (2024)). Instead of an iterative refinement process, these models learn to map noise to an image via a more direct, straight-line trajectory, representing a different theoretical foundation for high-quality generative modeling.
To mitigate the high computational cost of these iterative models, various acceleration techniques have been proposed. Model quantization (Li et al. (2024a; b; 2025c); Yan et al. (2025b)), cache mechanism (Xu et al. (2025); Pan et al. (2025)), sparse attention (Li et al. (2025a)), pruning (Wang et al. (2025a),Yan et al. (2025a)), and distillation (Hinton et al. (2015)) are general acceleration techniques for deep learning model. In diffusion models, specifically, one primary category is model quantization (Li et al. (2025b)), which reduces memory footprint and computational load by converting full-precision models into lower-bit representations. Another category involves cache mechanisms (Liu et al. (2025); Xu et al. (2018)), which enhance inference efficiency by exploiting temporal redundancy. These methods reuse intermediate features computed at earlier denoising steps to avoid redundant calculations in later steps. While effective in isolation, recent work like QuantCache (Wu et al. (2025)) demonstrates a unified framework can yield greater gains.
@@ 2.2 Editing Models
The task of editing real images with pre-trained generative models introduces the fundamental challenge of inversion: finding a latent representation that can faithfully reconstruct a given source image. This problem was first extensively studied in the context of Generative Adversarial Networks (GAN) Inversion (Wang et al. (2022),Zhu et al. (2020),Zhu et al. (2016)). In comparison, DDIM Inversion (Song et al. (2020b)) provides a deterministic method to find a corresponding noise latent for a source image. Once this latent is obtained, various editing mechanisms are employed during the denoising process to apply the desired changes. A prominent family of methods focuses on attention control, where the cross-attention maps between text and image are manipulated. For example, to change a “photo of a red car” to a “blue car,” Prompt-to-Prompt (Hertz et al. (2022)) identifies the attention weights corresponding to the word “red” and replaces them with those for “blue,” preserving the attention for “car” and the background. Another powerful technique is feature injection, exemplified by Plug-and-Play (PnP) (Zhang et al. (2021)). To preserve the identity of a subject, PnP injects the self-attention features—which encode structure and appearance—from the source image’s generation process into the edited one. A third approach is mask-based editing, where methods like DiffEdit (Couairon et al. (2022)) generate a mask indicating the region to be altered and then apply the denoising process only within that area. Despite these advances, a core challenge persists in perfectly disentangling the edited foreground from the unedited background.
@@ 3.1 One-Step Inversion-and-Editing
Motivation. Our motivation is to design a training strategy that explicitly decouples and progressively solves these two constraints. We posit that the network must first learn to jointly satisfy both objectives in a fully-supervised setting before it can be adapted to handle the unsupervised nature of real-image inversion. This leads to our ”Anchor-and-Refine” approach. The ”Anchor” stage uses synthetic data to ground the network in a parameter space that respects both constraints. The ”Refine” stage then adapts this mapping to real images, where we introduce a powerful generative prior from a teacher model to act as a proxy for the now-unsupervised distributional constraint. This ensures that even for real images, fidelity is pursued without sacrificing editability.
Proposed Method. Shown in Figure 2, our primary architectural modification is designed to resolve a fundamental tension in the inversion process. The inverted noise vector is typically burdened with two conflicting tasks: perfectly preserving the source image’s identity and remaining generic enough for subsequent editing. To decouple these roles, we introduce a dedicated visual adapter which provides the decoder $D$ with a direct visual information from the source image.
@@
$\mathcal{L}_{\text{Stage1}}=\lambda_{rec}\|z_{0}-\hat{z}_{0}\|_{2}^{2}+\lambda_{reg}\|\varepsilon_{gt}-\varepsilon_{inv}\|_{2}^{2}.$ (1)
The regression term $\mathcal{L}_{reg}$ is critical in this stage. It constrains the network to a region of the loss landscape where its outputs naturally conform to the target distribution $\mathcal{N}(0,I)$. During this stage, we train both the inversion network $I_{\theta}$ and the newly introduced image adapter. This teaches the adapter how to effectively provide visual priors that aid in reconstruction. This anchoring step prevents the network from converging to trivial solutions in the next stage.
@@ 3.2 Background Shield
Mask-Driven Recomposition and Selective Querying. During the editing pass with the target prompt $c_{tgt}$, we compute new queries, keys, and values ($Q_{tgt},K_{tgt},V_{tgt}\in\mathbb{R}^{S\times d_{k}}$). We then construct a spatially-aware, full key-value set, $K_{full},V_{full}$, by combining the background memory with the current foreground features:
$K_{full}[j,:]=\begin{cases}K^{*}_{\mathcal{B}}[\text{rank}_{\mathcal{B}}(j),:]&\text{if }j\in\mathcal{B}\\ K_{tgt}[j,:]&\text{if }j\in\mathcal{F}\end{cases},\quad V_{full}[j,:]=\begin{cases}V^{*}_{\mathcal{B}}[\text{rank}_{\mathcal{B}}(j),:]&\text{if }j\in\mathcal{B}\\ V_{tgt}[j,:]&\text{if }j\in\mathcal{F}\end{cases},$ (5)
where $\text{rank}_{\mathcal{B}}(j)$ ensures correct positional alignment. To mitigate boundary artifacts, we introduce a foreground core by applying a morphological erosion to the mask $M$. This is implemented using a 2D max-pooling operation (with kernel size $k$, stride $s$, and padding $p$) on the inverted mask. The resulting core mask $M_{core}$ is binarized with a threshold $\tau$ to yield the core index set $\mathcal{F}_{c}\subset\mathcal{F}$:
$M_{core}=\mathbf{1}-\text{MaxPool2d}(\mathbf{1}-M,kernel{\_}size,stride,padding),$ (6)
$\mathcal{F}_{c}=\{i\mid(M_{core})_{i}>\tau\}.$ (7)
The attention computation is then performed only for queries within this core region. Let $Q_{tgt,c}=Q_{tgt}[\mathcal{F}_{c},:]$ be the subset of queries corresponding to the core indices. The attention output for this region, $H_{c}\in\mathbb{R}^{|\mathcal{F}_{c}|\times d_{k}}$, is computed as:
$H_{c}=\text{softmax}\left(\frac{Q_{tgt,c}K_{full}^{T}}{\sqrt{d_{k}}}\right)V_{full}.$ (8)
The full output matrix $H\in\mathbb{R}^{S\times d_{k}}$ is then constructed by scattering the computed values $H_{c}$ back to their original positions, while all other positions corresponding to the background and boundary are set to zero.
Residual Fusion. The sparse output matrix $H$ is projected and added back to the input features: $Y=\text{Proj}(H)+X$. Since $H_{i}=0$ for all $i\notin\mathcal{F}_{c}$, this step functions as an identity map for the background and boundary regions, ensuring they are perfectly preserved.
@@ 3.3 Sparsified Spatial Cross-Attention
Identifying Key Semantic Tokens. Before computing attention, we first identify the most relevant tokens from the text prompt $y$ for the given edit region $M$. We compute the similarity between the set of image queries within the mask, $Q_{l,M}$, and all text keys $K_{y}$. The top-k text key-value pairs that exhibit the highest aggregate similarity are selected. This pre-selection step acts as a filter, creating a task-relevant subset of textual information, denoted as $(K_{y}^{k},V_{y}^{k})$.
@@
where $\text{rank}_{\mathcal{F}}(i)$ maps the global index to its local index within the foreground. Finally, this purified text guidance is integrated with the source image condition, $A_{img}$, to compute the updated hidden state $h_{l}$:
$h_{l}=s_{y}\cdot A_{\text{SSCA}}+s_{edit}\cdot M\odot A_{img}+s_{non-edit}\cdot(1-M)\odot A_{img}.$ (11)
This multi-step process provides a maximally disentangled and precise guidance signal for the edit.
Method Background Preservation CLIP Similarity
[…]
DDIM P2P 17.87 208.80 219.88 71.14 25.01 22.44
NT-Inv P2P 27.03 60.67 35.86 84.11 24.75 21.86
DDIM MasaCtrl 22.17 106.62 86.97 79.67 23.96 21.16
Direct Inversion MasaCtrl 22.64 87.94 81.09 81.33 24.38 21.35
DDIM P2P-Zero 20.44 172.22 144.12 74.67 22.80 20.54
Direct Inversion P2P-Zero 21.53 138.98 127.32 77.05 23.31 21.05
DDIM PnP 22.28 113.46 83.64 79.05 25.41 22.55
Direct Inversion PnP 22.46 106.06 80.45 79.68 25.41 22.62
ReNoise(SDXL) 20.85 176.84 51.78 72.44 24.41 21.88
TurboEdit 22.51 107.27 9.32 80.09 25.49 21.82
[…]
4 Experiment
4.1 Experimental Setup
Implementation Details. Our inversion network, $I_{\theta}$, is initialized from SwiftBrush (Nguyen & Tran (2024),Dao et al. (2024)). Inspired by (Song et al. (2024),Ye et al. (2023),Zhang et al. (2023)), the image conditioning branch is based on an adapter, utilizing a pre-trained CLIP image encoder. We train the model using the Adam optimizer (Kingma & Ba (2014)) with a learning rate of 2e-5, weight decay of 2e-4, and an Exponential Moving Average (EMA). Anchoring the Solution via Supervised Training runs for 150k iterationson synthetic data from SwiftBrush. Refining with a Teacher-Student Objective continues for 200k iterations using real images from CommonCanvas (Gokaslan et al. (2024)). All experiments were conducted on a single NVIDIA A6000 GPU.
Metrics. We evaluate our method on the PieBench benchmark (Zhang et al. (2021)), which features 700 samples across 10 editing types. We report metrics along two primary axes. As for Background Preservation, We compute PSNR (Huynh-Thu & Ghanbari (2008)), LPIPS (Zhang et al. (2018)), MSE and SSIM (Wang et al. (2004)) on the unedited regions to measure fidelity to the source image. As for Semantic Alignment, We report CLIP-Whole (Radford et al. (2021)) for prompt-image alignment and CLIP-Edited (Radford et al. (2021)) for alignment within the masked edit region.
Baselines. We compare our method against state-of-the-art multi-step and few-step baselines. For multi-step methods, we evaluate Prompt-to-Prompt (P2P) (Hertz et al. (2022)), MasaCtrl (Cao et al. (2023)), Pix2Pix-Zero (Parmar et al. (2023)), and Plug-and-Play (PnP) (Zhang et al. (2021)), paired with powerful inversion techniques like DDIM (Song et al. (2020a)), Null-text Inversion (NT-Inv) (Mokady et al. (2023a)), and Direct Inversion (Ju et al. (2023)). For few-step methods, we compare against Renoise (Garibi et al. (2024)) and TurboEdit (Deutch et al. (2024)).
@@ 4.1 Experimental Setup
- - 23.33 92.37 6.60 79.97 24.14 21.23
- 24.63 75.36 5.01 81.65 24.77 21.22
25.29 62.55 4.36 83.21 25.43 22.13

### v4

# FlashEdit: Decoupling Speed, Structure, and Semantics for Precise Image Editing

@@ Abstract
Figure 1: FlashEdit produces superior visual results for text-guided image editing, addressing background instability and semantic entanglement with an over $150\times$ speedup against DDIM [song2020denoising] + P2P [p2p].
footnotetext: * Equal contribution.footnotetext: † Corresponding author: Yulun Zhang, yulun100@gmail.com
@@ 1 Introduction
Text-guided image editing with diffusion models [instructpix2pix, dong2023prompt] has demonstrated remarkable capabilities, allowing users to perform complex semantic modifications with high fidelity. The standard methodology is built upon a two-stage inversion-denoising pipeline: an initial inversion process maps a source image to its corresponding noise latent, which is then progressively denoised to generate the edited output according to a target prompt [directInversion, masactrl]. The objective is to achieve high fidelity in both content preservation and target prompt alignment, which often necessitates a computationally intensive, multi-step process.
Recent research has pursued several distinct strategies to improve accuracy and speed. To tackle the latency of the multi-step denoising process, methods based on model distillation have been proposed to enable editing in a faster way [turboedit]. These approaches must carefully address challenges such as mismatched noise statistics and insufficient editing strength that arise when adapting multi-step frameworks to fast samplers [mokady2023null, miyake2025negative]. To improve edit precision and prevent semantic leakage into the background, another category of work modifies the model’s internal mechanisms, primarily by re-weighting or replacing attention maps to ensure the edit is spatially constrained [alphaedit, headrouter]. Recognizing that the final edit quality is highly dependent on the starting point, other approaches focus on refining the inversion technique itself [directInversion]. These methods aim to find a more accurate initial latent vector, with recent insights revealing that separating the objectives of content preservation and edit fidelity can yield significant performance gains and speedups [wang2025image].
@@
  • At the temporal level, we introduce the One-Step Inversion-and-Editing (OSIE) pipeline and its “Anchor-and-Refine” training strategy, which enables high-fidelity inversion for one-step diffusion models.
  • At the spatial level, we propose Background Shield (BG-Shield), a structural intervention in self-attention that uses memory caching and soft recomposition with attention feathering to enforce pixel-perfect background preservation, ensuring the structural integrity of the edit.
@@ 2.1 Diffusion Models
Recent advances in image synthesis have been largely driven by diffusion models [peebles2023scalable, kulikov2024flowedit], which have become a leading paradigm for generating high-fidelity images from text. The core mechanism involves an iterative denoising process that progressively refines a random noise vector into a coherent image conditioned on a text prompt. A landmark contribution in this area is Stable Diffusion  [stableDiffusion], a Latent Diffusion Model (LDM) [rombach2022high] that performs the computationally intensive denoising process in a lower-dimensional latent space, making the technology widely accessible. Parallel to this, alternative frameworks have emerged, such as Flow Matching models like Flux [flux2024]. Instead of an iterative refinement process, these models learn to map noise to an image via a more direct, straight-line trajectory, representing a different theoretical foundation for high-quality generative modeling.
To mitigate the high computational cost of these iterative models, various acceleration techniques have been proposed. Model quantization [li2024large, li2024fast, arbllm, pbs2p], cache mechanism [specee, specdiff], sparse attention [sparseAttn], pruning [specprunevla, yan2025recalkv], and distillation [hinton2015distilling, nguyen2025swiftedit] are general acceleration techniques for deep learning model. In diffusion models, specifically, one primary category is model quantization [dvdquant], which reduces memory footprint and computational load by converting full-precision model weights and activations into lower-bit representations. Another category involves cache mechanisms [liu2025reusing, deepcache, TaylorSeer2025], which enhance inference efficiency by exploiting temporal redundancy. These methods reuse intermediate features computed at earlier denoising steps to avoid redundant calculations in later steps. While effective in isolation, recent work like QuantCache [quantcache] demonstrates a unified framework can yield greater gains.
@@ 2.2 Editing Models
The task of editing real images with pre-trained generative models introduces the fundamental challenge of inversion: finding a latent representation that can faithfully reconstruct a given source image. This problem was first extensively studied in the context of Generative Adversarial Networks (GAN) Inversion [wang2021HFGI, zhu2020domain, zhu2016generative]. In comparison, DDIM Inversion [song2020denoising] provides a deterministic method to find a corresponding noise latent for a source image. Once this latent is obtained, various editing mechanisms are employed during the denoising process to apply the desired changes. A prominent family of methods focuses on attention control, where the cross-attention maps between text and image are manipulated. For example, to change a “photo of a red car” to a “blue car,” Prompt-to-Prompt [p2p] identifies the attention weights corresponding to the word “red” and replaces them with those for “blue,” preserving the attention for “car” and the background. Another powerful technique is feature injection, exemplified by Plug-and-Play (PnP) [pnp]. To preserve the identity of a subject, PnP injects the self-attention features—which encode structure and appearance—from the source image’s generation process into the edited one. A third approach is mask-based editing, where methods like DiffEdit [couairon2022diffedit] generate a mask indicating the region to be altered and then apply the denoising process only within that area. Despite these advances, a core challenge persists in perfectly disentangling the edited foreground from the unedited background.
@@ 3.1 One-Step Inversion-and-Editing
Motivation.  Our motivation is to design a training strategy that explicitly decouples and progressively solves these two constraints. We posit that the network must first learn to jointly satisfy both objectives in a fully-supervised setting before it can be adapted to handle the unsupervised nature of real-image inversion. This leads to our “Anchor-and-Refine” approach. The “Anchor” stage uses synthetic data to ground the network in a parameter space that respects both constraints. The ”Refine” stage then adapts this mapping to real images, where we introduce a powerful generative prior from a teacher model to act as a proxy for the now-unsupervised distributional constraint. This ensures that even for real images, fidelity is pursued without sacrificing editability.
Proposed Method.  Shown in Figure 2, our primary architectural modification is designed to resolve a fundamental tension in the inversion process. The inverted noise vector is typically burdened with two conflicting tasks: perfectly preserving the source image’s identity and remaining generic enough for subsequent editing. To decouple these roles, we introduce a dedicated visual adapter which provides the decoder $D$ with a direct visual information from the source image, inspired by [yin2024one, nguyen2025swiftedit, chen2025dove].
@@
$\mathcal{L}_{\text{Stage1}}=\|z_{0}-\hat{z}_{0}\|_{2}^{2}+\|\varepsilon_{gt}-\varepsilon_{inv}\|_{2}^{2}.$ (1)
The $\|\varepsilon_{gt}-\varepsilon_{inv}\|_{2}^{2}$ term is critical in this stage. It constrains the network to a region of the loss landscape where its outputs naturally conform to the target distribution $\mathcal{N}(0,I)$. During this stage, we train both the inversion network $I_{\theta}$ and the image adapter. This teaches the adapter how to effectively provide visual priors that aid in reconstruction. This anchoring step prevents the network from converging to trivial solutions in the next stage.
@@
In addition, we introduce a Cycle-Consistency Regularization loss, $\mathcal{L}_{\text{cycle}}$, to enforce that our inversion network $I_{\theta}$ and the one-step generator $G$ are functional inverses. We first compute the initial noise $\varepsilon_{inv}=I_{\theta}(z_{0},c_{src})$ and its reconstruction $\hat{z}_{0}=G(\varepsilon_{inv},c_{src})$. We then feed this reconstruction back into the inversion network to get a cycled noise, $\varepsilon_{cycle}=I_{\theta}(\hat{z}_{0},c_{src})$. The loss is defined as:
$\mathcal{L}_{\text{cycle}}=\left\|\varepsilon_{inv}-\varepsilon_{cycle}\right\|_{2}^{2}$ (4)
This loss ensures the noise from the real image is consistent with the noise from its own reconstruction.
@@ 3.2 Background Shield
Soft Recomposition with Attention Feathering.  During the editing pass with the target prompt $c_{tgt}$, we compute new queries, keys, and values ($Q_{tgt},K_{tgt},V_{tgt}\in\mathbb{R}^{S\times d_{k}}$). We then construct a spatially-aware, full key-value set, $K_{full},V_{full}$, by combining the background memory with the current foreground features:
$K_{full}[j,:]=\begin{cases}K^{*}_{\mathcal{B}}[\text{rank}_{\mathcal{B}}(j),:]&\text{if }j\in\mathcal{B}\\ K_{tgt}[j,:]&\text{if }j\in\mathcal{F}\end{cases}\quad$ (6)
$V_{full}[j,:]=\begin{cases}V^{*}_{\mathcal{B}}[\text{rank}_{\mathcal{B}}(j),:]&\text{if }j\in\mathcal{B}\\ V_{tgt}[j,:]&\text{if }j\in\mathcal{F}\end{cases}$ (7)
where $\text{rank}_{\mathcal{B}}(j)$ ensures correct positional alignment.
To mitigate boundary artifacts and ensure a smooth blend, we replace the hard-masking of the original “foreground core” approach with an Attention Feathering mechanism. We first create a soft alpha mask $A\in[0,1]^{S}$ by applying a Gaussian blur to the binary mask $M$:
$A=\text{GaussianBlur}(M,\text{kernel\_size}=\sigma)$ (8)
This alpha mask creates a “transition zone”. This zone softly connects the editing foreground, where the mask value is close to one, and the preserved background, where the mask value is close to zero.
Next, we compute the full attention output, $H_{edit}\in\mathbb{R}^{S\times d_{k}}$, where all queries $Q_{tgt}$ attend to the recomposed key-value set $(K_{full},V_{full})$:
$H_{edit}=\text{softmax}\left(\frac{Q_{tgt}K_{full}^{T}}{\sqrt{d_{k}}}\right)V_{full}.$ (9)
The final sparse output matrix $H$ is then obtained by $H_{edit}$ with our soft alpha mask $A$:
$H=A\odot H_{edit}$ (10)
where $\odot$ denotes element-wise multiplication.
Method Background Preservation CLIP Similarity
[…]
DDIM [ddim] P2P [p2p] 17.87 208.80 219.88 71.14 25.01 22.44
NT-Inv [NT-Inv] P2P [p2p] 27.03 60.67 35.86 84.11 24.75 21.86
DDIM [ddim] MasaCtrl [masactrl] 22.17 106.62 86.97 79.67 23.96 21.16
Direct Inversion [directInversion] MasaCtrl [masactrl] 22.64 87.94 81.09 81.33 24.38 21.35
DDIM [ddim] P2P-Zero [pix2pixZero] 20.44 172.22 144.12 74.67 22.80 20.54
Direct Inversion [directInversion] P2P-Zero [pix2pixZero] 21.53 138.98 127.32 77.05 23.31 21.05
DDIM [ddim] PnP [pnp] 22.28 113.46 83.64 79.05 25.41 22.55
Direct Inversion [directInversion] PnP [pnp] 22.46 106.06 80.45 79.68 25.41 22.62
ReNoise(SDXL) [renoise] 20.85 176.84 51.78 72.44 24.41 21.88
TurboEdit [turboedit] 22.51 107.27 9.32 80.09 25.49 21.82
SwiftEdit [nguyen2025swiftedit] 23.31 71.54 6.18 82.25 25.56 21.91
[…]
@@ 3.3 Sparsified Spatial Cross-Attention
Identifying Key Semantic Tokens.   Before computing attention, we first identify the most relevant tokens from the text prompt $y$ for the given edit region $M$. We compute the similarity between the set of image queries within the mask, $Q_{l,M}$, and all text keys $K_{y}$. The top-k text key-value pairs that exhibit the highest aggregate similarity are selected. This pre-selection step is a filter, creating a task-relevant subset of textual information, denoted as $(K_{y}^{k},V_{y}^{k})$.
@@
where $\text{rank}_{\mathcal{F}}(i)$ maps the global index to its local index within the foreground. Sparsified Spatial Cross-Attention (SSCA) provides a maximally disentangled and precise guidance signal for the edit.
@@
- - 23.38 92.29 6.58 80.01 24.17 21.28
- 24.69 75.28 4.96 81.95 24.80 21.28
25.33 62.25 4.28 83.32 25.51 22.29
4 Experiment
4.1 Experimental Settings
Implementation Details.   We implement our framework and architecture in PyTorch inspired by [wang2025osdface, nguyen2025swiftedit, chen2025dove]. The image conditioning branch is inspired by recent approaches [moma, ip-adapter, controlnet]. We train the model using the Adam optimizer [adam] with standard hyperparameters. The first stage, Anchoring the Solution via Supervised Training, is run on synthetic data [sun2023journeydb]. The second stage, Refining with a Teacher-Student Objective, uses real images [commoncanvas]. Experiments were conducted on a NVIDIA A6000 GPU.
Metrics.   We evaluate our method on the PieBench benchmark [pnp], which features 700 samples across 10 editing types. We report metrics along two primary axes. As for Background Preservation, We compute PSNR [PSNR], LPIPS [lpips], MSE and SSIM [ssim] on the unedited regions to measure fidelity to the source image. As for Semantic Alignment, We report CLIP-Whole [clip] for prompt-image alignment and CLIP-Edited [clip] for alignment within the masked edit region.
Baselines.   We compare our method against state-of-the-art multi-steps and few-steps baselines. For multi-step methods, we evaluate Prompt-to-Prompt (P2P) [p2p], MasaCtrl [masactrl], Pix2Pix-Zero [pix2pixZero], and Plug-and-Play (PnP) [pnp], paired with powerful inversion techniques like DDIM [ddim], Null-text Inversion (NT-Inv) [NT-Inv], and Direct Inversion [directInversion]. For few-steps methods, we compare against Renoise [renoise], TurboEdit [turboedit] and SwiftEdit [nguyen2025swiftedit].